OpenStudy (anonymous):

which of the following is a solution of the equation 5x + 3y = -7? a. (1,-2) b. (-1,2) c. (-2,1) d. (2,-1)

OpenStudy (shubhamsrg):

why dont you try and satisfy these points on the eqn yourself?

OpenStudy (anonymous):

i dont know how?

OpenStudy (anonymous):

could you teach me?

OpenStudy (shubhamsrg):

(a,b) means a is x co-co-ordinate and b is y co-ordinate..helpful now ?

OpenStudy (anonymous):

wowowowowowowo,,,,,,,,, how old are u ?

OpenStudy (anonymous):

am about 7 y.o.

OpenStudy (kaederfds):

the answer is c. (-2,1) just put the answer choices into the question

OpenStudy (anonymous):

shubhamsrg: still not, could you explain more

OpenStudy (shubhamsrg):

it'll take a lot of time since you maybe dont know the basics yet..guess you should google over,,or consult wikipedia..

OpenStudy (anonymous):

5x + 3y = -7,,,, for the answer is (X,Y) so if any value with (-2,1) it's mean for x=-2 and for y=1 and then put the x and y value to u'r equation,,,,,, so we yield : 5.(-2) + 3.(1) = -7 -10+3= ...........??? can u complete this kirscute ,,,,,,,???